Solve Cos x + Sin y = Cos 2x + Sin 2x [ x=2n pi x= 2n pi/3 + pi/6]

Solve Cos x  + Sin y = Cos 2x + Sin 2x [ x=2n pi x= 2n pi/3 + pi/6]

Grade:11

1 Answers

Chirag jain
5 years ago
FIRSTLY IN PLACE OF SINY THERE WILL BE SINX
Cosx-cos2x=sin2x-sinx
2sin(3x/2)sin(x/2)=2sin(3x/2)cos(x/2)w
Sin3x{sin(x/2)-cos(x/2)}=0
Now,case1=}sin(3x/2)=sin0
                       3x/2=npi+(-1)^n×0
                       x=2npi/3_____1
        Case2=} sin(x/2)=cos(x/2)
                        tan(x/2)=tan(pi/4)
                        x/2=npi+pi/4
                        x=2npi+pi/2    or, 
                        x=2npi/3+pi/6
Thnx to give me a chance